OpenStudy (anonymous):

Find a polynomial of minimum degree (there are many) that has the given zeros -7, 07,

hartnn (hartnn):

i assume thats -7,0,7 if a is zero, x-a is a factor. hence the three factors here are x(x-7)(x+7) did u get this ?

hartnn (hartnn):

so then using, (x-7)(x+7)=x^2-49 your polynomial will be f(x)=x^3-49x ok?

OpenStudy (anonymous):

your answer is correct, thanks :D

hartnn (hartnn):

welcome :)

OpenStudy (anonymous):

heres another one

hartnn (hartnn):

new question, new post.